# FLoC: Facility Location-Based Efficient Visual Token Compression for Long Video Understanding

### v1

@@ 4.2 Benchmarks
Video-MME  Fu et al. (2024) is a multi-modal evaluation benchmark designed to assess various dimensions of visual and textual understanding in videos. It includes diverse real-life footage covering different domains such as sports, news, and user-generated content. The dataset focuses on tasks like video captioning, event detection, and question answering. Its complexity stems from the highly varied video lengths and content types, challenging models to handle both short clips and longer sequences effectively.
LongVideoBench.  Wu et al. (2025) is explicitly curated for the study of long-form video understanding, featuring videos that run significantly longer than those found in typical benchmarks. These videos cover a range of categories—lectures, live events, and continuous surveillance footage—where crucial information may appear sporadically or in non-consecutive segments. The benchmark focuses on tasks such as topic segmentation, extended event detection, and global summarization.
MLVU (Multi-Level Video Understanding) Zhou et al. (2024) is a dataset aiming to evaluate hierarchical comprehension of videos, from low-level frame recognition to high-level storyline interpretation. It comprises clips sourced from various genres, including movies, documentaries, and instructional videos.   [^2] VideoMME, LongvideoBench, and MLVU are all under the CC BY-SA 4.0 International License.
@@ 4.2.1 Implementation
We effectively evaluated the performance of various visual token compression algorithms using the lmms-eval toolkit Li et al. (2024); Zhang et al. (2024b) as our codebase, which supports multiple video LMM models and diverse benchmarks. All experiments were conducted leveraging NVIDIA H100 GPUs and multiprocessing for efficient computation.
@@ 4.3 Baselines
Recent Algorithms We compared the performance of recently proposed algorithms, LongVU Shen et al. (2024), DyCoke Tao et al. (2024), TS-LLaVA Qu et al. (2024), PruneVID Huang et al. (2024), and DivPrune Alvar et al. (2025). Implementation details are described in Section E of Appendix.
Model Qwen2.5-VL-7B
Max
Frames
Method
Video
MME
MLVU LVB Avg.
768 - 66.33 70.31 60.51 65.72
7200 TS-LLaVA 65.07 72.40 62.08 66.52
LongVU 65.04 71.02 62.75 66.27
DivPrune 64.93 70.19 62.30 65.81
Random 64.56 70.52 61.63 65.57
DyCoke 65.78 71.30 62.98 66.69
PruneVid 62.96 68.63 62.45 64.68
[…]
Model Qwen2.5-VL-32B
Max
Frames
Method
Video
MME
MLVU LVB Avg.
768 - 70.41 71.57 62.60 68.19
7200 TS-LLaVA 70.22 73.09 65.00 69.44
LongVU 70.37 72.22 64.62 69.07
DivPrune 70.26 73.37 64.32 69.32
Random 69.70 72.49 64.62 68.94
DyCoke 71.00 72.26 63.87 69.04
PruneVid 68.00 70.19 63.50 67.23
[…]
@@ 4.4 Results
Table 1 presents a comparative analysis of video understanding performance using Video-MME, MLVU, and LongVideoBench (LVB) with 6 different baseline methods as described in 4.3. We evaluate these methods under various visual token compression ratios of $2^{-5}$, $2^{-4}$, and $2^{-3}$. Figure 1 illustrates the performance of each compression algorithm in terms of accuracy retention (x-axis), measured as a percentage relative to the full-token baseline, and compression time (y-axis). The results are based on a 1/8 compression ratio using the Qwen2.5-VL-7B model.
As shown in the results, our method consistently outperforms existing visual token compression techniques across different datasets, compression ratios, and backbone models. While comparing algorithms generally improve the video understanding performances by incorporating structured token selection, they may still overlook less frequent but contextually significant elements, which our method successfully retains. DivPrune can also select diverse visual token sets based on min-max diversity problem, but it does not consider the coverage of visual tokens.
@@
As shown in Table 2, FLoC can significantly improve the performance of LMMs that are conventionally measured using a limited number of frames. For the 7B model, the accuracy increased by an average of 1.31 points, and for the 32B model, it rose by an average of 2.44 points. These results indicate that while existing LMMs are forced to process fewer frames due to their limited context length, our proposed algorithm enables them to handle a larger number of frames through efficient compression. We believe this approach substantially enhances their overall video understanding capabilities.
@@ 4.5.1 Representative and Diverse Visual Tokens
Additionally, as shown in Fig. 5, our proposed FLoC selects diverse tokens, successfully capturing visual cues like hats and sunglasses, unlike DivPrune and TS-LLaVA, which often miss them. This enables more accurate answers to questions about what the woman is wearing. Additional results with more examples are provided in the Appendix, specifically illustrated in Figure 7 and  8.
@@ 4.5.2 Minimal Computational Overheads
Table 4: Robustness of our proposed method across diverse ranges of block lengths, denoted as $T$.
Ratio $T=1$ $T=2$ $T=4$ $T=8$ $T=16$ $T=32$ $T=64$ $T=256$
$2^{-3}$ 59.33 60.37 61.04 61.48 61.56 60.89 60.81 60.26
$2^{-4}$ 57.70 58.07 59.04 59.11 59.11 59.41 58.63 58.93
$2^{-5}$ 54.93 54.59 56.52 56.30 57.00 57.70 57.85 56.93
@@ 4.5.3 Robustness on Block Lengths
Table 4 illustrates the performance across a wide range of block lengths, denoted as $T$ which is the only hyper-parameter for our method. Here, We use Qwen-VL 7B and ViodeMME for the model and dataset, respectively. Our method demonstrates consistent and robust performance across various $T$ values. Despite its robustness, using an excessively small $T$ may fail to capture temporally similar visual tokens, while a significantly large $T$ introduces substantial computational overhead. Thus, we selected an optimal block length of $T=32$ excepting for the result of $T=8$ in Figure 1.
@@ 5 Conclusion
As long video understanding advances, handling the overwhelming number of visual tokens remains a key bottleneck. While prior methods such as uniform sampling and clustering have addressed this issue, they often fail to capture sufficient visual diversity and add computational overhead. We tackle these limitations by proposing a visual token compression framework based on the facility location function. Our method selects tokens that are both representative and diverse, preserving essential scene information while significantly reducing computation via a lazy greedy algorithm. Extensive experiments on large-scale benchmarks, including Video-MME, LongVideoBench, and MLVU show that our method consistently outperforms existing compression techniques. Its efficiency and strong performance without added overhead make it well-suited for real-world applications such as surveillance, augmented reality, and autonomous navigation. As video-LMMs scale, improving efficiency and information retention will be key to advancing long video understanding.
@@ Appendix E Baseline Implementation Details
  • PruneVID: We utilized the query-agnostic compression part involving spatial-temporal token merging stage. The implementation was based on the authors’ official Github repository.
@@
Appendix G Limitations and Future Directions

### v2

@@ 4.2 Benchmarks
Video-MME (Fu et al., 2024) is a multi-modal evaluation benchmark designed to assess visual and textual understanding in videos, covering diverse real-life footage across domains such as sports, news, and user-generated content. It focuses on tasks like video captioning, event detection, and question answering.
LongVideoBench (Wu et al., 2025) is curated for long-form video understanding, featuring extended videos such as lectures, live events, and surveillance footage, emphasizing topic segmentation and global summarization.
MLVU (Multi-Level Video Understanding) (Zhou et al., 2024) evaluates hierarchical comprehension from frame-level recognition to storyline interpretation, using clips from movies, documentaries, and instructional videos. [^2] Video-MME, LongVideoBench, and MLVU are all under the CC BY-SA 4.0 International License.
EgoSchema (Mangalam et al., 2023) evaluates egocentric video understanding through short, first-person perspective clips, emphasizing schema-level reasoning and activity prediction.
Among these, Video-MME, LongVideoBench, and MLVU include videos longer than one hour, making them suitable for long-form video understanding. In contrast, EgoSchema consists of relatively short, minute-level clips but remain widely adopted benchmarks for video understanding research.
@@ 4.2.1 Implementation
We effectively evaluated the performance of various visual token compression algorithms using the lmms-eval toolkit (Li et al., 2024; Zhang et al., 2024b) as our codebase, which supports multiple video LMM models and diverse benchmarks. All experiments were conducted leveraging NVIDIA H100 GPUs and multiprocessing for efficient computation.
@@ 4.3 Baselines
Model Qwen2.5-VL-7B Qwen2.5-VL-32B
Max Frames Method Video MME MLVU LVB Avg. Video MME MLVU LVB Avg.
768 - 66.33 70.31 60.51 65.82 70.41 71.57 62.60 68.19
7200 TS-LLaVA 65.07 72.40 62.08 66.52 70.22 73.09 65.00 69.44
LongVU 65.04 71.02 62.75 66.27 70.37 72.22 64.62 69.07
DivPrune 64.93 70.19 62.30 65.81 70.26 73.37 64.32 69.32
Random 64.56 70.52 61.63 65.57 69.70 72.49 64.62 68.94
DyCoke 65.78 71.30 62.98 66.69 71.00 72.26 63.87 69.04
PruneVid 62.96 68.63 62.45 64.68 68.00 70.19 63.50 67.23
[…]
Recent Algorithms We compared the performance of recently proposed algorithms, LongVU (Shen et al., 2024), DyCoke (Tao et al., 2024), TS-LLaVA (Qu et al., 2024), PruneVID (Huang et al., 2024), DivPrune (Alvar et al., 2025), STTM (Hyun et al., 2025), LLaVA Scissor (Sun et al., 2025), and FastVID (Shen et al., 2025). Implementation details are described in Section E of Appendix.
@@ 4.4 Results
Table 1 presents a comparative analysis of video understanding performance using above-mentioned 6 benchmarks with 9 different baseline methods as described in 4.3. We evaluate these methods under various visual token compression ratios of $2^{-5}$, $2^{-4}$, and $2^{-3}$. Figure 1 illustrates the performance of each compression algorithm in terms of accuracy retention (x-axis), measured as a percentage relative to the full-token baseline, and compression time (y-axis). The results are based on a 1/8 compression ratio using the Qwen2.5-VL-7B model.
As shown in the results, our method consistently outperforms existing visual token compression techniques across different datasets, compression ratios, and backbone models. We attribute this superiority to FLoC’s ability to overcome the structural limitations of prior approaches. Specifically, graph-based merging methods (e.g., STTM, LLaVA Scissor) often suffer from the “weak connection” problem, where distinct tokens—such as small objects and their background—are irreversibly merged based on local similarity thresholds, leading to significant detail loss especially at low compression ratios. Similarly, while diversity-based methods (e.g., DivPrune) effectively capture outliers, they often fail to retain representative tokens that describe the core context of the video. In contrast, our facility location-based approach mathematically guarantees a balance between representativeness and diversity, successfully retaining both the central narrative and fine-grained visual cues that other methods overlook.
@@
As shown in Table 2, FLoC can significantly improve the performance of LMMs that are conventionally measured using a limited number of frames. For the 7B model, the accuracy increased by an average of 1.21 points, and for the 32B model, it rose by an average of 2.44 points. These results indicate that while existing LMMs are forced to process fewer frames due to their limited context length, our proposed algorithm enables them to handle a larger number of frames through efficient compression. We believe this approach substantially enhances their overall video understanding capabilities.
@@ 4.5.1 Representative and Diverse Visual Tokens
Additionally, as shown in Fig. 5, our proposed FLoC selects diverse tokens, successfully capturing visual cues like hats and sunglasses, unlike DivPrune and TS-LLaVA, which often miss them. This enables more accurate answers to questions about what the woman is wearing. Additional results with more examples are provided in the Appendix, specifically illustrated in Figure 8 and  9.
@@ 4.5.3 Robustness on Block Lengths
Figure 6: Performance versus block length $T$ (log2-scale) for a number of benchmark datasets.
To examine the impact of the sole hyperparameter in our proposed algorithm—the block length TT—we evaluated performance across various datasets and compression ratios while varying $T$. In this experiment, the Qwen2VL-7B model was used.
As illustrated in Figure 6, we observe distinct behaviors depending on the block length $T$. In the region where $T\leq 4$, performance tends to degrade because the narrow temporal window prevents the algorithm from identifying redundancy across adjacent blocks (inter-block redundancy). Conversely, as $T$ increases, the facility location objective optimizes representativeness and diversity over a broader temporal context, leading to performance saturation. Crucially, unlike traditional clustering methods where computational cost scales quadratically with input size, our lazy greedy implementation ensures that increasing $T$ incurs negligible latency overhead. This suggests that a sufficiently large fixed block length (e.g., $T=32$) serves as a robust and efficient default, minimizing the need for per-video hyperparameter tuning. However, we acknowledge that relying on fixed uniform segmentation is a heuristic simplification and may not be strictly optimal for every video content. We anticipate that developing an adaptive mechanism to automatically determine $T$ based on temporal dynamics could yield further performance improvements. A more detailed discussion on this limitation and potential future directions is provided in Appendix I
@@ 5 Conclusion
As long video understanding advances, handling the overwhelming number of visual tokens remains a key bottleneck. While prior methods such as uniform sampling and clustering have addressed this issue, they often fail to capture sufficient visual diversity and add computational overhead. We tackle these limitations by proposing a visual token compression framework based on the facility location function. Our method selects tokens that are both representative and diverse, preserving essential scene information while significantly reducing computation via a lazy greedy algorithm. Extensive experiments on large-scale benchmarks, including Video-MME, LongVideoBench, MLVU, and Egoschema show that our method consistently outperforms existing compression techniques. Its efficiency and strong performance without added overhead make it well-suited for real-world applications such as surveillance, augmented reality, and autonomous navigation. As video-LMMs scale, improving efficiency and information retention will be key to advancing long video understanding.
@@ Appendix E Baseline Implementation Details
  • PruneVID: We utilized only the first stage of the algorithm, which performs query-agnostic spatial-temporal token merging. To ensure a fair comparison, the subsequent query-aware stage was excluded. All experiments were conducted based on the official GitHub repository provided by the authors.
@@
  • STTM and LLaVA Scissor: We used the official implementations provided by the authors for all benchmarks. Hyperparameters were kept at their default settings, while threshold values were adjusted to achieve the desired compression ratio.
  • FastVID: We conducted experiments based on the official GitHub repository provided by the authors. Among the models we tested, implementation was available only for the Qwen2.5-VL model; therefore, experiments were limited to this model. We varied only the retention ratio while keeping all other hyperparameters at their default values.
@@
Appendix G Profiling of Computational and Memory Footprint
Method Inference (s) Compression (s) Total (s) GFLOPS VRAM (GB)
Full 3.22 2.69 27.33
FLoC 1.70 0.99 3.00 318 17.96
DyCoke 0.13 1.83 0.46 17.96
PruneVID 0.19 1.89 1.69 17.96
STTM 0.07 1.77 0.17 17.96
Scissor 1.24 2.94 783 17.96
FastVID 0.44 2.14 6 17.96
DivPrune 2.37 4.07 317 17.96
LongVU 0.38 2.08 317 17.96
kmeans 12.31 14.01 82 17.96
kmedoids 2.41 4.11 82 17.96
spectral 8.12 9.82 82 17.96
Table 7: Performance comparison of different methods
To compare the resource consumption and speed of our proposed algorithm against baseline methods, we measured LLM inference time, compression time, FLOPs, and peak VRAM usage. All experiments were conducted using the Qwen2.5-VL 7B model on an NVIDIA H100 GPU, with a compression ratio of 12.5%. The reported times correspond to end-to-end processing of 784 frames, each containing 60 visual tokens. FLOPs were estimated by accumulating the actual operation counts during execution, with approximations for modularized components (which may introduce minor inaccuracies). Peak VRAM usage was recorded using PyTorch’s torch.cuda.max_memory_allocated() function.
Overall, graph-based methods tend to achieve faster compression and require fewer operations, but—as shown in previous experiments—they exhibit inferior performance. In contrast, clustering-based methods incur significantly higher computational costs, resulting in slower processing speeds. Our proposed algorithm performs compression in less time than the LLM inference step, demonstrating practical efficiency. However, it exhibits relatively high FLOPs, most of which are attributed to pairwise similarity computations.
Appendix H Ablation Study on Similarity Metrics in Facility Location
To investigate the impact of the similarity measure on the facility location function, we conducted an ablation study comparing our default Cosine Similarity with Euclidean Distance. Since the facility location function requires a similarity matrix, we converted the Euclidean distance into a similarity measure using a Gaussian kernel:
$S_{euc}(x,y)=\exp\left(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\right)$
where $\sigma$ is set to the median of all pairwise Euclidean distances within the set, following the standard median heuristic. We performed experiments using the InternVL3-8B model across different compression ratios (1/8, 1/16, and 1/32).
Table 8: Performance with different distance metrics.
$T$ Metric VideoMME MLVU LVB EgoSchema Average Difference
$2^{-3}$ Cosine 64.93 71.57 56.69 69.40 65.65 +0.55
Euclidean 64.26 72.26 58.49 69.80 66.20
$2^{-4}$ Cosine 63.41 69.09 56.47 66.20 63.79 -0.27
Euclidean 62.74 69.50 55.42 66.40 63.52
$2^{-5}$ Cosine 60.81 66.93 54.23 63.80 61.44 -0.97
Euclidean 59.59 65.59 53.70 63.00 60.47
The results are summarized in Table 8. We observed that while the Euclidean-based metric showed a slight advantage (+0.55 accuracy) at a low compression ratio (1/8), Cosine Similarity consistently outperformed Euclidean similarity as the compression ratio increased. Specifically, Cosine similarity achieved higher accuracy at 1/16 (+0.27) and 1/32 (+0.97) ratios.
This suggests that while Euclidean distance captures fine-grained magnitude differences useful when retaining many tokens, Cosine similarity is more robust for abstract feature space coverage, particularly in high-compression regimes where capturing the dominant semantic directions is crucial. Based on these findings, we adopted Cosine similarity as the default metric to ensure consistent performance across varying degrees of compression.
Appendix I Limitations and Future Directions